estimate the number of molecules in 3.00m^3 of air by assuming that it has a density of 1.21kg/m^3 and that is entirely composed of nitrogen molecules
1
Expert's answer
2020-02-18T10:16:51-0500

Total volume of the gas=3.0"m^3"

Density of the molecule ="1.21 kg\/m^3"

we know that the atomic weight of the nitrogen molecule="28.014 gm\/mol"

We know that,

mass = density x volume

"m= 3\times 1.21 kg"

"m=3.63kg"

Hence the number moles="\dfrac{3.63\times 1000}{28.014}=129" .578

We know that the avagadro number (number of molecules per mole)="6.023\times 10^{23}"


Hence the number of nitrogen molecules "=129.578\times 6.023\times 10^{23}=780.45\times 10^{23}" molecules.